OpenStudy (anonymous):

The graph of y=3x^2+12x+19 is a parabola that opens in the direction______. It has vertex____,y-intercept____, and ____real roots.

hero (hero):

positive direction. It has vertex -b/2a, y intercept at x = 0, and 2 real roots

OpenStudy (anonymous):

vertex = -12/3 =-4

OpenStudy (anonymous):

no, vertex = -2

OpenStudy (anonymous):

upward direction (-2,7) y=19 2 real roots